### v1

@@ 5.3 Regression Performance
Besides predicting the risk of adverse outcomes, we also examined the applications of 3D body shape in estimating current intrauterine parameters. We adjusted the final layer of our network and conducted regression experiments to estimate fetal weight and MVP. In these experiments, we designed a similar baseline method using Support Vector Regression (SVR) with traditional anthropometric measurements as inputs. We also included the average of the ground truth values as an additional baseline. Linear regression (LR), BPNN, RF, and SVR with PCA were tested as comparisons methods , using the same inputs as in the classification tasks. The evaluation was conducted using 5-fold cross-validation, consistent with the approach in the classification experiments. The results are presented in Table 8 and Table 8 below.
@@
Baseline 34.31 47.57 9.33% 13.36% 62.79% 39.53%
Average 52.04 71.25 13.71% 17.93% 55.81% 20.93%
PCA+LR 41.54 56.42 11.39% 15.73% 55.81% 39.53%
PCA+BPNN 39.40 55.84 10.94% 16.05% 67.44% 34.88%
PCA+RF 43.26 60.69 11.64% 16.60% 62.79% 41.86%
PCA+SVR 38.44 48.27 10.46% 13.62% 60.46% 30.23%
Ours 30.41 43.62 8.28% 12.31% 76.74% 53.49%
@@
Baseline 0.786 0.988 16.18% 20.20% 37.5% 22.5%
Average 0.747 0.929 15.68% 19.85% 42.50% 27.50%
PCA+LR 1.248 1.552 25.02% 29.79% 20.00% 15.00%
PCA+BPNN 1.126 1.848 23.57% 29.32% 27.50% 12.50%
PCA+RF 0.838 1.091 17.60% 22.97% 45.00% 27.50%
PCA+SVR 0.830 1.035 17.12% 21.16% 40.00% 20.00%
Ours 0.976 1.137 20.29% 24.28% 17.50% 12.50%
Figure 6: Regression results of EFW.
The results from Table 8 indicate that our algorithm outperforms other algorithms in estimating EFW across all metrics. While SVR also demonstrates commendable performance, our algorithm exceeds SVR by 20.89% in MAE, 9.63% in RMSE, 20.84% in MAPE, 9.62% in RMSPE, and achieves an accuracy rate of 76.74% within a 10% error tolerance range. Interestingly, the baseline method also yields promising results, outperforming the same SVR algorithm with 3D body shape as inputs. Figure 6 illustrates the estimation result of each participant calculated by different algorithms. For clarity, the participants were reordered by increasing order of their actual EFW values collected from ultrasound report. From Figure 6 we find that except the average line, all other estimation results follow the trend of the actual value, with our method yielding results that fluctuate the least around the ground truth. These findings align with the observations from Table 8, establishing the utility of 3D body shape in estimating fetal weight.
When we attempted to estimate MVP value using these algorithms, the performances were not as good as when estimating fetal weights, as shown in Table 8. The MAPEs and RMSPEs of machine learning methods on this task exceed 15% and 20% respectively, which is notably higher than the roughly 10% MAPEs and 15% RMSPEs when estimating fetal weight. The RF algorithm achieves the highest accuracies, but none of these machine learning methods surpass the estimation using average value of the ground truth. This suggests that 3D body shape may not be an effective tool for estimating MVP. The line chart with reordered subjects is provided in Figure 7. In contrast to the results in Figure 6, we can barely find the relevance among the distributions of these estimation results, and none appears to align with the increasing trend of the actual value.
Figure 7: Regression results of MVP.
@@ 5.4 Ablation Study
Our proposed method exhibits overall good performance in both classification and regression tasks, which may be attributed to the separate application of PCA and RNN to extract complementary features. To substantiate this claim, we conducted a set of experiments that estimate fetal weight using networks with or without either of these two streams. Additionally, we explored replacing the RNN with the gated recurrent unit (GRU), an enhanced version of the vanilla RNN, to extract sequential information from level circumferences. The GRU is designed with ability of learning both long-term and short-term dependencies, while having fewer parameters than the long short-term memory (LSTM) network. Through comparisons with this GRU network, we aim to demonstrate that our network architecture performs better with the conventional RNN. Table 9 lists the results of the various combinations we tested during the algorithm development process.
@@
(PCA+RNN+FC) 30.41 43.62 8.28% 12.31% 76.74%
Comparing the results between the second and third rows in Table 9, we find that PCA facilitates FC in learning more useful information from raw level circumferences, leading to a 4.28g decrease in MAE and 18.60% increase in accuracy. Further comparison between the third, fourth and last rows indicates that using either PCA or RNN alone does not yield optimal results. The incorporation of both, as implemented in our final design, presents the most significant improvement in performance. The fifth row shows the performance results when the RNN is replaced with GRU in our network. Despite the GRU’s advanced capabilities in capturing long-term relationships from sequences, this modification does not bring more improvement in this task. This failure could potentially be due to the fact that GRU contains more trainable parameters than the conventional RNN, thus resulting in reduced learning efficiency, particularly when dealing with small data size. These findings affirm the effectiveness of our proposed hybrid structure for 3D body feature extraction.
@@ 6 Conclusion
In this study, we examined the potential of using 3D body scans with few basic demographic information for prenatal care for future telehealth application and explored how to use this modality efficiently. We proposed a novel neural network which incorporates a supervised learning stream and an unsupervised learning stream for extracting features from sampled abdominal level circumferences. The supervised stream leverages RNN units to extract sequential information from level circumferences, while the unsupervised stream employs the PCA to captures global descriptors of abdominal body shape. We applied this algorithm to predict the risk of preterm labor, GDM, GH, the likelihood of cesarean section, and to estimate current fetal weight and MVP. The results indicate that maternal 3D body shape, captured during 18-24 gestational weeks, is effective in predicting GDM, GH, and preterm labor, as well as estimating fetal weight. In comparison to other well-performing machine learning algorithms, our proposed algorithm demonstrates superior performances in these tasks.
Nevertheless, there is room for future improvement. Firstly, our current dataset is small and only contains scans obtained in patients’ second trimester of pregnancy, potentially limiting the model’s generalizability to estimate fetal weight in other trimesters, since fetal growth trends may vary across different stages. Ideally, with the inclusion of actual birth weights and 3D body scans closer to delivery, a more accurate model may be developed guided by neonatal weights. We may also explore the potential of using longitudinal data for better estimation. Secondly, our current analysis focused solely on the abdominal region, overlooking other body parts. Given the research [17, 47, 48, 49] suggesting correlations between appendicular body shape and pregnancy outcomes, future studies could investigate methods using the entire body shape to get more accurate estimations and predictions. To enable the practical application of 3D body scans in telehealth, we will also conduct further experiments using smartphone-scanned 3D models.

### v2

@@ 5.3 Regression Performance
In addition to predicting the risk of adverse outcomes, we also explored the use of 3D body shape for estimating current intrauterine parameters. We adjusted the final layer of our network and conducted regression experiments to estimate fetal weight and MVP. A baseline method was designed using a Linear Regression (LR) model with traditional anthropometric measurements as inputs, and we also included the mean of the ground truth values as an additional baseline. LR, BPNN, RF, and SVR with PCA were evaluated using the same inputs as in the classification tasks for comparison. The results on the test sets are presented in Table 8 and Table 8 below. The best results are also highlighted in bold.
@@
Baseline 34.80 44.53 9.43% 11.57% 61.11% 30.56%
Average 43.53 62.03 11.20% 14.53% 55.56% 30.56%
PCA+LR 39.25 51.88 10.51% 13.63% 58.33% 30.56%
PCA+BPNN 36.03 50.24 9.62% 13.03% 61.11% 38.89%
PCA+RF 33.54 45.09 9.08% 11.84% 61.11% 41.67%
PCA+SVR 34.12 44.37 9.07% 11.39% 61.11% 30.56%
Ours 30.40 42.78 8.19% 10.94% 72.22% 44.44%
@@
Baseline 0.718 0.880 16.80% 21.69% 35.29% 23.53%
Average 0.542 0.727 12.21% 16.22% 50.00% 41.18%
PCA+LR 0.921 1.075 21.20% 25.72% 26.47% 11.76%
PCA+BPNN 0.696 0.873 16.19% 21.28% 44.12% 20.59%
PCA+RF 0.672 0.848 15.57% 20.35% 35.29% 29.41%
PCA+SVR 0.672 0.851 15.80% 20.67% 50.00% 20.59%
Ours 0.625 0.801 14.50% 19.21% 50.00% 29.41%
Figure 6: Bland-Altman plot of regression results for EFW.
The results in Table 8 indicate that our algorithm outperforms other algorithms in estimating EFW across all metrics. While the SVR model also demonstrates commendable performance, our algorithm exceeds it by 10.90% in MAE, 3.58% in RMSE, 9.70% in MAPE, 3.95% in RMSPE, and achieves an accuracy of 72.22% within a 10% error tolerance. Interestingly, the baseline method also shows promising results, outperforming the same LR model using 3D body shape as inputs. Figure 6 illustrates the estimation results for each participant using different algorithms with Bland-Altman plots, which are widely used to analyze the agreement between two clinical measurements in biomedicine. In these plots, the mean of the differences is shown as a red dash-dotted line, and the 95% limits of agreement are represented by two dashed lines. From Figure 6, we find that the estimations of our method distribute more tightly within a narrower agreement interval. This finding aligns with the observations from Table 8, demonstrating the utility of 3D body shape in estimating fetal weight. We also investigated the mutual outlier with an average value around 410 and find that its EFW percentile among all fetuses at the same GA is 99%. It indicates that body shape–based methods still have room for improvement in extreme cases. Moreover, by observing the potential fitted line of the scatter plots, we find that all of these methods tend to underestimate the lower fetal weights and overestimate the higher fetal weights, which can be addressed in future work.
When estimating MVP using these algorithms, the performances are not as good as when estimating fetal weights, as shown in Table 8. For machine learning methods other than ours, the MAPEs and RMSPEs on this task exceeded 15% and 20% respectively, which is notably higher than the roughly 9% MAPEs and 13% RMSPEs for fetal weight estimation. Although our method achieves the highest accuracy, none of the machine learning methods outperform the simple baseline using the mean of the ground truth. This suggests that 3D body shape may not be an effective predictor of MVP. The corresponding Bland-Altman plots are shown in Figure 7. From these plots, we find that none of the machine learning methods achieve a narrower agreement interval than the mean-value prediction. Moreover, based on the fitted lines of the scatter plots, the three methods listed in the bottom row of Figure 7 tend to predict values close to the average rather than capturing meaningful patterns. This unexpected result may be due to fluctuations caused by factors such as fetal position and maternal hydration [50, 51]. Additionally, MVP is less related to maternal nutritional status, which is a key feature captured by body shape–based methods.
Figure 7: Bland-Altman plot of regression results for MVP.
@@ 5.4 Ablation Study
Our proposed method exhibits overall good performance in both classification and regression tasks, which may be attributed to the separate application of PCA and RNN to extract complementary features. To substantiate this claim, we conducted a set of experiments that estimate fetal weight using networks with or without either of these two streams. Additionally, we explored replacing the RNN with the gated recurrent unit (GRU), an enhanced version of the vanilla RNN, to extract sequential information from level circumferences. The GRU is designed with ability of learning both long-term and short-term dependencies, while having fewer parameters than the long short-term memory (LSTM) network. Through comparisons with this GRU network, we aim to demonstrate that our network architecture performs better with the conventional RNN. Table 9 lists the results of the various combinations we tested during the algorithm development process with the best performances highlighted in bold.
@@
(PCA+RNN+FC) 30.40 42.78 8.19% 10.94% 72.22%
Comparing the results between the second and third rows in Table 9, we find that PCA facilitates FC in learning more useful information from raw level circumferences, leading to a 1.03g decrease in MAE and 15.78% increase in accuracy. Further comparison between the third, fourth and last rows indicates that using either PCA or RNN alone does not yield optimal results. The incorporation of both, as implemented in our final design, presents the most significant improvement in performance. The fifth row shows the performance results when the RNN is replaced by GRU in our network. Despite the GRU’s advanced capabilities in capturing long-term relationships from sequences, this modification does not bring more improvement in this task. This failure could potentially be due to the fact that GRU contains more trainable parameters than the conventional RNN, thus resulting in reduced learning efficiency, particularly when dealing with small data size. These findings affirm the effectiveness of our proposed hybrid structure for 3D body feature extraction.
@@ 6 Conclusion
In this study, we investigated the potential of using 3D body scans with basic body shape-related demographic information for prenatal care in future telehealth applications and explored how to use this modality effectively. We proposed a novel neural network which incorporates a supervised learning stream and an unsupervised learning stream for extracting features from sampled abdominal level circumferences. The supervised stream leverages RNN units to extract sequential information from level circumferences, while the unsupervised stream employs the PCA to captures global descriptors of abdominal body shape. We applied this algorithm to predict the risk of preterm labor, GDM, PEC, the likelihood of cesarean section, and to estimate current fetal weight and MVP. The results indicate that maternal 3D body shape, captured during 18-24 gestational weeks, is effective in predicting GDM and preterm labor, as well as estimating fetal weight. Compared to other well-performing machine learning algorithms, our proposed method demonstrates superior performances in these tasks.
Nevertheless, there is room for future improvement. Firstly, our current dataset is small and only contains scans obtained in patients’ second trimester of pregnancy, potentially limiting the model’s generalizability to estimate fetal weight in other trimesters, since fetal growth trends may vary across different stages. Ideally, with the inclusion of actual birth weights and 3D body scans closer to delivery, a more accurate model may be developed guided by neonatal weights. We may also explore the potential of using longitudinal data for better estimation. Secondly, our current analysis focused solely on the abdominal region, overlooking other body parts. Given the research [17, 52, 53, 54] suggesting correlations between appendicular body shape and pregnancy outcomes, future studies could investigate methods using the entire body shape to get more accurate estimations and predictions. Thirdly, this pilot study only explored the usability of body shape for the discussed tasks, however, additional easily accessible medical variables, such as obstetric history and chronic disease history, could be incorporated to increase model precision. With these variables incorporated, it can be more practically applicable for intelligent, remote prenatal care. Additionally, to further facilitate the application of 3D body scans in telehealth, we will also conduct extended experiments using 3D models obtained from smartphone scans.